OpenStudy (anonymous):

Subtract 6 1/2 - 4 7/10 . Justify your steps.

OpenStudy (anonymous):

change into improper fractions \[\frac{ 13 }{ 2 }-\frac{ 47 }{ 10 }\]

OpenStudy (anonymous):

find a common denominator\[\frac{ 13 }{ 2 }\times \frac{ 5 }{ 5 }=\frac{ 65 }{ 10 }\]

OpenStudy (anonymous):

then subtract\[\frac{ 65 }{ 10 }-\frac{ 47 }{ 10 }=\frac{ 18 }{ 10 }\]

OpenStudy (anonymous):

now simplify

OpenStudy (anonymous):

reduce

OpenStudy (anonymous):

thankyou so much :)

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

should the fraction be 1 4/5 before i reduce

OpenStudy (phi):

the final answer can be 18/10 (not reduced, but still valid , unless they ask for reduced) divide top and bottom by 2 to get 9/5 (improper fraction) or 1 4/5 (same form as the problem, so a good choice) another variation is 18/10 is 1.8 (same as 1 4/5, but in decimal form) for this problem, where you started with mixed numbers, I would put the answer in mixed number form, which is what you did.

OpenStudy (anonymous):

thankyou i will become a fan

OpenStudy (anonymous):

one rule of thumb in math is always to reduce fractions even if the problem doesn't ask you to. If your teacher looks at the problem his/her answer will be in reduces form and he/she will greatly appreciate that yours is as well. :)